# Generalized Onsager’s Relation in Magnon Hall Effect and Its Implication

### v2

@@ Abstract
The two Onsager’s relations in Eq. (10) and (Generalized Onsager’s Relation in Magnon Hall Effect and Its Implication) are main results of this work. They reveal complicated dependence of the magnon Berry curvature on different types of exchange couplings. The first Onsager’s relation can obviously reproduce the usual understanding that the magnon Hall effect is induced by the DM interaction with vanishing anisotropic parameter $\Gamma_{ij}$. Besides that, the magnon Hall effect can have much richer structures.
Implications.—One direct consequence of the second Onsager’s relation is that the isotropic exchange interaction can be as important as the DM interaction in determining the magnon Hall effect. In fact, the role of the isotropic exchange and that of the DM interaction are even interchangeable by a proper gauge transformation. On one hand, with vanishing anisotropic exchange coupling, the two Onsager’s relations suggest that the magnon Hall conductivity is an odd function of both the DM interaction and the isotropic exchange coupling. On the other hand, from Eq. (Generalized Onsager’s Relation in Magnon Hall Effect and Its Implication), an additional $\pi/2$ phase factor in $A_{ij}$ inverts its real and imaginary part and hence leads to $J_{ni,mj}^{+}\rightarrow-D_{ni,mj}$ and $D_{ni,mj}\rightarrow J_{ni,mj}^{+}$. Physically, such a phase factor amounts to rotating the spin frame at either one of the $i$ and $j$ sites around the local $z$ axis by $\pi/2$, which should not generate any measurable difference.
Figure 2: (a) Honeycomb lattice with in-plane magnetization pointing $45^{\circ}$ from the horizontal line. (b) The thermal conductivity as a function of $J_{N}$ and $D$. In (b), $J_{A}=0.1\sqrt{J_{N}^{2}+D^{2}}$, $J_{B}=0.15\sqrt{J_{N}^{2}+D^{2}}$; the blue and red curve have positive and negative sign, respectively.
As a concrete example, we consider a ferromagnetic honeycomb lattice with nearest and next nearest exchange coupling, DM interaction and easy-plane anisotropy. The corresponding magnetic Hamiltonian reads as
$H=$ $-\sum_{\langle i,j\rangle}{J_{N}\bm{S}_{i}\cdot\bm{S}_{j}}-\sum_{\langle\langle i,j\rangle\rangle}{\sum_{\mu=A,B}{J_{\mu}\bm{S}_{i}^{\mu}\cdot\bm{S}_{j}^{\mu}}}$ (12)
$+\sum_{\langle i,j\rangle}{D\left(\bm{R}_{ij}\times\hat{\bm{z}}\right)\cdot\left(\bm{S}_{i}\times\bm{S}_{j}\right)}.$
where $\bm{R}_{ij}$ is the unit vector from site $i$ to $j$. In the nearest neighbor exchange coupling, when $J_{A}=J_{B}$, the chiral symmetry $\mathcal{C}=\sigma_{x}K$ is respected, and the system is hence gapless with vanishing magnon Hall effect. Moreover, an external magnetic field is implied which stabilizes the in-plane ferromagnetic order, as shown in Fig. 2(a).
In Fig. 2(b), we plot the Hall conductivity as a function of $J_{N}$ and $D$, with $J_{N}^{2}+D^{2}$ fixed. We find that $\kappa_{xy}$ has four lobes with alternating signs, resembling the angular dependence of $d_{xy}$ wave. As a result, $\kappa_{xy}$ is indeed an odd function of $J_{N}$ and $D$, consistent with the two Onsager’s relations. Such sign-change of the magnon Hall conductivity can potentially be observed in materials with ferromagnetic-antiferromagnetic transition.
The $d$-wave pattern further suggests the equivalence between $J_{N}$ and $D$. To see this, we note that each A site has three nearest-neighbor bonds, which we label by $1$, $2$, $3$, as shown in Fig. 2. The exchanging coupling parameter on each bond forms a two-dimensional vector, with $(J_{1},D_{1})=(J_{N},-D/\sqrt{2})$, $(J_{2},D_{2})=(J_{N},D\cos(\pi/12))$, and $(J_{3},D_{3})=(J_{N},-D\sin(\pi/12))$. A rotation of the local spin frame at site A then interchanges the role of $J_{i}$ and $D_{i}$: $(J_{i},D_{i})\rightarrow(D_{i},-J_{i})$ with $i=1,2,3$. More specifically, since $\kappa_{xy}$ should be unchanged under such rotation of local spin frame, only the relative orientation angles between the three $(J_{i},D_{i})$ survive. The leading order contribution to $\kappa_{xy}$ should then has the following pattern: $\kappa_{xy}=A_{1}(J_{1},D_{1})\times(J_{2},D_{2})+A_{2}(J_{1},D_{1})\times(J_{3},D_{3})+A_{3}(J_{2},D_{2})\times(J_{3},D_{3})$, where $A_{i}$ is a function of the magnitude of the three vectors $(J_{\ell},D_{\ell})$, and the cross product of two two-dimensional vector generates a pseudoscalar. Therefore, $\kappa_{xy}$ is a function of $J_{N}\cdot D$, consistent with the d-wave pattern and clearly demonstrates the equivalence between $J_{N}$ and $D$. Moreover, since $|D_{i}|<D$, $J_{i}^{2}+D_{i}^{2}$ is always slimmer in the $D$-direction, as can be observed in Fig. 2(b).
Another consequence of the two Onsager’s relations is that the DM interaction is not necessary for the magnon Hall effect. It can be fully replaced by the anisotropic exchange coupling. By setting $D=0$, Eq. (10) suggests that $\kappa$ is an odd function of the anisotropic exchange parameter $\Gamma$. Moreover, similar to the equivalence between $D$ and $J^{+}$, for the anisotropic exchange coupling, the off-diagonal part $\Gamma_{ni,mj}$ is equivalent to $J_{ni,mj}^{-}$, due to a rotation of the local spin frame at either $i$ or $j$ by $\pi/2$.
Figure 3: (a) The thermal conductivity as a function of $\Gamma$ and $J^{-}$. Parameters are $J^{+}=\pm 0.5J_{0}$ which changes sign as $\Gamma$ varies, $\sqrt{(J^{-})^{2}+\Gamma^{2}}=0.1J_{0}$, $J_{A}=0.05J_{0}$, and $J_{B}=0.075J_{0}$. The blue and red curve are positive and negative, respectively. (b) Symmetry requirements for DM interaction(top) and anisotropic exchange coupling(bottom).
To show this, we consider the following ferromagnetic honeycomb model with the magnetic order being along $z$-axis:
$H=\sum_{\langle i,j\rangle}{\bm{S}_{i}^{T}\mathcal{J}_{ij}\bm{S}_{j}}-\sum_{\langle\langle i,j\rangle\rangle}{\sum_{\mu=A,B}{J_{\mu}\bm{S}_{i}^{\mu}\cdot\bm{S}_{j}^{\mu}}}\,,$ (13)
where
$\mathcal{J}_{ij}=\left(\begin{matrix}-J^{+}+J^{-}&w_{ij}\Gamma&0\\ w_{ij}\Gamma&-J^{+}-J^{-}&0\\ 0&0&-J_{0}\\ \end{matrix}\right).$ (14)
where $\Gamma$, $J^{+}$ and $J^{-}$ are constants representing off-diagonal anisotropic, isotropic, diagonal anisotropic exchange couplings, respectively. $w_{ij}$ are weighing factors for the three nearest-neighbor bonds, and they are chosen to break the in-plane mirror symmetry. The next nearest neighbor term is added to break the chiral symmetry as before.
In Fig. 3 we plot the thermal Hall conductivity as a function of $\Gamma$ and $J^{-}$. One can find a similar $d$-wave pattern, showing that $\kappa_{xy}$ is an odd function of both $\Gamma$ and $J^{-}$, and the equivalence between $\Gamma$ and $J^{-}$. We shall emphasize that the anisotropic exchange coupling only affects $J^{-}$ and $\Gamma$ in our case, while the one in previous study[7] also serves as antisymmetric exchange interaction in the magnon Hamiltonian[35]. Therefore, our work unambiguously demonstrates that the anisotropic exchange alone can support the magnon Hall effect.
The anisotropic exchange coupling has completely different symmetry requirements from the DM interaction. The latter requires the breaking of inversion symmetry. In comparison, the former requires the breaking of mirror-x and mirror-y symmetry; the inversion symmetry can be retained. This then extends the material candidate for the magnon Hall effect.
@@ Acknowledgements.
We acknowledge useful discussions with Shiyue Deng, Shu Li and Qingtao Zhao. This work is supported by the National Natural Science Foundation of China (12234017). Y. G. is also supported by the Innovation Program for Quantum Science and Technology (2021ZD0302802). The supercomputing service of USTC is gratefully acknowledged.

### v3

@@ Abstract
The role of the symmetric exchange interaction.—The two Onsager’s relations in Eq. (7) and (Magnon Thermal Hall Effect Induced By Symmetric Exchange Interaction) are main results of this work. They reveal complicated dependence of the magnon Berry curvature on different types of exchange couplings. The first Onsager’s relation can obviously reproduce the usual understanding that the magnon thermal Hall effect is induced by the DM interaction with vanishing anisotropic parameter $\Gamma_{ij}$.
More importantly, Eq. (7) suggests that when the DM intereaction is zero, the magnon thermal Hall effect can still emerge, purely due to the symmetric but anisotropic exchange interaction. In this case, $\kappa_{\mu\nu}$ is an odd function of $\Gamma_{ij}$. Equation (Magnon Thermal Hall Effect Induced By Symmetric Exchange Interaction) further yields a necessary condition for this case: for the symmetric exchange between two different sites $i$ and $j$, the $J_{ij}^{\pm}$ is also necessary and the resulting magnon thermal Hall effect is also an odd function of $J_{ij}^{\pm}$. In fact, one can further prove that $\Gamma_{ni,mj}$ is equivalent to $J_{ni,mj}^{-}$, due to a rotation of the local spin frame at either $i$ or $j$ by $\pi/2$.
Figure 1: DM interaction (a) and symmetric exchange interaction (b). In (a), the intermediate non-magnetic gray atom breaks the inversion and mirror-z symmetry. In (b), the mirror-x and mirror-y symmetry should be broken to have the symmetric coupling $S_{ix}S_{jy}+S_{iy}S_{jx}$. The mirror-z symmetry is allowed.
The symmetric exchange interaction has different symmetry requirements from the DM interaction. The latter requires the breaking of local inversion symmetry. In contrast, to have nonzero symmetric exchange interaction, the local inversion symmetry can exist, and only certain local mirror symmetry needs to be broken, as shown in Fig.1. For example, to have $J_{xy}(S_{ix}S_{jy}+S_{iy}S_{jx})$, mirror-x and mirror-y symmetry should be broken and mirror-z symmetry can be retained. We shall note that to make this term affect the magnon thermal Hall effect, the equilibrium spin direction should contain perpendicular component to the retained mirror plane.
Previously, it is known that besides the DM interaction, the dipole-dipole interaction can also cause a magnon thermal Hall effect[5, 43, 44]. Here our theory of the symmetric exchange interaction includes the dipole-dipole interaction as a special case. In fact, the latter has the following form $\hat{H}_{dipole}\propto(\hat{e}_{ij}\cdot\bm{S}_{i})(\hat{e}_{ij}\cdot\bm{S}_{j})$, which is always symmetric upon interchanging $\bm{S}_{i}$ with $\bm{S}_{j}$. However, since the dipole-dipole interaction comes from the electromagnetic force, it is usually much smaller than the one from the electron-electron exchange interaction, provided that the required symmetry breaking for the latter is achieved.
To illustrate this type of magnon thermal Hall effect, we consider a honeycomb lattice with in-plane antiferromagnetic order, as shown in Fig. 2(a). We choose the honeycomb lattice as there are two magnetic atoms in a unit cell, which is essential for generating a nontrivial magnon Berry curvature. Otherwise, one should consider a sample with finite dimension along one direction to achieve the same purpose. For each nearest neighbor bond, we assume that the inversion symmetry and the mirror symmetry that interchanges neighboring magnetic atoms are retained. Therefore, the nearest neighbor DM interaction vanishes. The Hamiltonian is given by
$H=\sum_{\langle i,j\rangle}{\bm{S}_{i}^{T}\mathcal{J}_{ij}\bm{S}_{j}}-\sum_{i}\bm{B}\cdot\bm{S}_{i}.$ (9)
The first term takes account the isotropic and anisotropic symmetric exchange coupling, consistent with the symmetry, and the second term is the Zeeman energy from external magnetic field along the $x$-direction, which breaks the chiral symmetry $\sigma_{x}K$ and hence induces a finite band gap for a nontrivial Berry curvature [40]. The detail of parameters can be found in Ref. [40].
In Fig. 2(b) we plot the thermal Hall conductivity as a function of $\Gamma$ and $J^{-}$. One can find a $d$-wave pattern, showing that $\kappa_{xy}$ is an odd function of both $\Gamma$ and $J^{-}$, consistent with the two Onsager’s relation in Eq. (7) and (Magnon Thermal Hall Effect Induced By Symmetric Exchange Interaction). It demonstrates the necessary condition for this type of thermal Hall conductivity, i.e., a nonzero $J^{-}$ is also required.
In the electronic anomalous Hall effect, it is often assumed that the magnetization direction is perpendicular to the Hall-deflection plane due to a widely used empirical law. However, in recent years, it has been found that if the crystal symmetry is sufficiently low, the magnetization can lie within the Hall-deflection plane, causing an in-plane anomalous Hall effect [45, 46, 47, 48, 49].
Figure 2: Magnon thermal Hall effect from symmetric exchange interaction. (a) Antiferromagnetic honeycomb lattice, with mirror planes bisecting each nearest neighbor bond. The spin order is along the x direction. (b) The thermal conductivity as a function of parameter $\alpha$ with $\alpha$ being the angle of $J^{-}+i\Gamma$. The magnitude $\sqrt{(J^{-})^{2}+\Gamma^{2}}$ is fixed. (c) Schematics of in-plane magnon thermal Hall effect. (d) Thermal conductivity as a function of the orientation of the in-plane ferromagnetic spin order aligned by external magnetic field.
Such in-plane geometry can persist in the magnonic degree of freedom with the help of the anisotropic exchange interaction. To show this, we consider the same honeycomb lattice as above but with in-plane ferromagnetic spin order. The spin order can then be aligned by external magnetic field, as shown in Fig. 2(c). The magnon Hamiltonian is similar with Eq. (9). An additional staggered isotropic exchange term between second nearest neighbors is added to break the chiral symmetry, which can be achieved by strain gradient, Zeeman field gradient, in-plane electric field, etc..
We then calculate the thermal Hall conductivity as a function of the orientation of the in-plane spin order. As shown in Fig. 2(d), the result is nonzero and exhibits a three-fold rotational symmetry, consistent with the $C_{3}$ rotational symmetry. Moreover, when the rotation angle is $\pi$, i.e., the spin order is flipped, the spin frame for magnons also change by a $C_{2x}$ rotation which reverses the sign of $\Gamma$ without changing other properties of the magnon spectrum. According to Onsager’s relation in Eq. (7), thermal conductivity should also flip sign, as shown in Fig. 2(d). This exotic pattern is analogous to those discovered in electronic anomalous Hall effect [50, 51, 52, 53].
Symmetrywise speaking, the generation of the in-plane magnon thermal Hall effect is similar to that of the in-plane anomalous Hall effect, i.e., there should be no $C_{2}$ axis perpendicular to the Hall-deflection plane[54]. The only out-of-plane rotation axis allowed is thus $C_{3}$. In-plane $C_{2}$ axis is also allowed, and when the spin order aligns with this axis, the thermal conductivity vanishes. This occurs in materials with point group $D_{3}$, $D_{3d}$.
We shall emphasize that the symmetry analysis is only for a specific Hall-deflection plane. As for the original crystals, the point group symmetry can be quite high. For example, for materials with cubic point group such as $T$, $T_{d}$, $T_{h}$, $O$, $O_{h}$, the in-plane magnon thermal Hall effect can still be discovered as long as the Hall deflection plane is properly selected to fulfill the above symmetry requirement, such as the $(112)$-plane.
The mechanism from the symmetric but anisotropic exchange interaction expands the potential systems with magnon thermal Hall effect, in the way that it allows re-examination of magnetic crystals that preserves local inversion symmetry or that the local DM interaction contains a $\bm{D}$ vector perpendicular to the spin order. One promising candidate is ferromagnetic material $\mathrm{VAu_{4}}$. The middle point of the nearest V atoms is an inversion center[40] so that DM interaction is forbidden. Moreover, the two mirror symmetries shown in FIG.2(b) are broken and hence the symmetric exchange interaction is allowed which can cause a nonzero magnon thermal Hall effect in a thin-film $\mathrm{VAu_{4}}$. Similar systems include layered $\mathrm{CrCl_{3}}$ and $\mathrm{V_{2}Se_{2}O}$ with in-plane electric field or strain gradient.
In summary, we establish two generalized Onsager’s relations in magnon thermal Hall effect from which we propose a distinct mechanism of the magnon thermal Hall effect solely from the symmetric but anisotropic exchange interaction. The mechanism can compete with the previously proposed mechanisms from the DM interaction and dipole-dipole interaction. More importantly, as it involves different types of symmetry breaking, it can expands systems with the magnon thermal Hall effect. As an application, we show that it can leads to a special in-plane magnon thermal Hall effect.
@@ Acknowledgements.
We acknowledge useful discussions with Zheng Liu, Shiyue Deng, Shu Li and Qingtao Zhao. This work is supported by the National Natural Science Foundation of China (12234017). Y. G. is also supported by the Innovation Program for Quantum Science and Technology (2021ZD0302802). The supercomputing service of USTC is gratefully acknowledged.